OpenStudy (anonymous):

factor the binomials completely. a^2 - b^4

OpenStudy (anonymous):

another difference of two squares\[a^2 - b^4=(a)^2-(b^2)^2=(a-b^2)(a+b^2)\]

OpenStudy (anonymous):

ty again

OpenStudy (anonymous):

np again :})